OpenStudy (anonymous):

how to transform (^6√(x^5))^7 into an expression with a rational exponent?

OpenStudy (kropot72):

\[\large \sqrt[6]{x^{5}}=x^{\frac{5}{6}}\] Therefore we now can write the original expression as \[\large (x^{\frac{5}{6}})^{7}\] Can you do the last step?

OpenStudy (kropot72):

@badsuns Are you there?

OpenStudy (kropot72):

You need to use the rule \[\large (x^{a})^{b}=x^{ab}\]

OpenStudy (anonymous):

sorry sorry, i was doing other work. i figured out the answer though. thanks anyway! :)